I experience an experiments 2 times (experiments A and B) now I want to know which one is better the other in percentage , find the ratio of them how to find how many percentage dose A enhanced my task I also want to know if what i did below is true?
Assume A=1.5541 and B=0.9531 in unit of second of execution time of two algorithm then:
A/B=1.630573917 ==> A is 1.6 longer than B
1-(B(ie. smaller value)/A))= 0.39 *100 ==>> B enhanced 39% in execution time

B/A*100=what we call this result is it true to say B is 61% better than A
I am working on Performance of computer system I was away from math and statistics could you give me some reference(usefull book) that can help me in calculations like these or even better one that can relate to problems like performance measurments?.
Thanks

Answer by stanbon(75887)

I experience an experiments 2 times (experiments A and B) now I want to know which one is better the other in percentage , find the ratio of them how to find how many percentage dose A enhanced my task I also want to know if what i did below is true?
Assume A=1.5541 and B=0.9531 in unit of second of execution time of two algorithm then:
A/B=1.630573917 ==> A is 1.6 longer than B
1-(B(ie. smaller value)/A))= 0.39 *100 ==>> B enhanced 39% in execution time
B/A*100=what we call this result is it true to say B is 61% better than A
I am working on Performance of computer system I was away from math and statistics could you give me some reference(usefull book) that can help me in calculations like these or even better one that can relate to problems like performance measurments?.
-------
A time - B time = 1.5541-0.9531 = 0.6010
---
Then 0.6010/1.5541 = 0.3867 = 38.67%
B is 38.67% faster than A
--------------------------------
Cheers,
Stan H.
===============